Home
Ask directory
Intro Stats / AP Statistics
April 2024
2
Intro Stats / AP Statistics
Q&A Archive of
April 2, 2024
Select Question
April 2 of 2024
1. A-F Suppose we observe x_(1),x_(2),dots,x_(n) as independent and identically distributed as an exponen- tial random variable with f_(lambda )(x_(1))=lambda e^(lambda _(1)), and recall that E(x_(1))=(1)/(lambda ) and Var(x_(1))=(1)/(lambda ^(2)). (a) Write down the conditional density…
For the given data set representing the runs scored by two players in the last 10 matches, conduct the following analysis: egin{center} egin{tabular}{|c|c|c|c|c|c|c|c|c|c|c|} hline Match & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 9 & 10 \ hline Player A & 8 & 42 & 56 & 68 & 91 & 123 & 12 & 46 & 57 &…
Problem 4 (15 points) A call center receives 25 callers per minute on average. On average, a caller spends 1 minute on hold and 3.5 minutes talking to a service representative. On average, how many callers are "in" the call center (meaning that they are either on hold or talking to a service…
A Box-and-Whisker plot for the sample data set "Some Data" is shown, along with some additional statistics about the data. ar{x} =48.9 s=16.5,n=83 a. Find each of the following: Range: Median: Interquartile Range: b. What proportion of the data falls between the values of 37 and 71 ? (Enter…
13.15 Suppose Y is a continuous random variable having the density function fY(y)=1/π(1+y^2) for y∈R. (a) Verify that this is a valid density function. (Hint: Use symmetry, and remember trig substitution!) (b) Try to find the expected value of Y and show that it does not exist. 13.17 Suppose…
Nico will perform a hypothesis test at the alpha =0.01 level. Evie will perform the same test at the alpha =0.10 level. Part: 0 / 2 Part 1 of 2 (a) If H_(0) is true, who has a greater probability of making a Type I error? If H_(0) is true, | has a greater probability of making a Type I…
From a binomial probability distribution with n=20, select any probability of success p between 0.1 and 0.4 from a cumulative binomial probability table and use the table to find: a) P(x>=5) and b) P(4<=x<=10) c) P(x=6) 3.From a binomial probability distribution with n-20, select any…
As N becomes smaller, x (standard error) becomes stays the same. larger. closer in value to the mean. smaller. As N becomes smaller,X (standard error) becomes stays the same. larger. closer in value to the mean. smaller.
5.5-2. Let x be N(75,64). Using the same set of axes, sketch the graphs of the probability density functions of (a) x. (b) ar{x} , the mean of a random sample of size 16 from this distribution. (c) ar{x} , the mean of a random sample of size 36 from this distribution. 5.5-2. Let X be N(75,…
(d) If an order comes in for 10,000 steel rods, how many rods should the plant manager expect to manufacture if the order states that all rods must be between 27.9 cm and 28.1 cm?
Predicting expected revenue in dollars from a new customer, predicting sales revenue for a store, and predicting the length of stay in a hospital based on patient characteristics and medical conditions are examples of: a. Applicability of statistical techniques b. Applicability of…
For a normal distribution, exactly 89.97% of the z-score values are less than z = +1.50. True False
For this exercise, you will be navigating problem 10.13, Sexually Transmitted Disease, and problem 10.15, Ophthalmology, both on page 439 of the textbook. Here are a few helpful hints: For problem 10.13, construct a 2x2 contingency table of the data presented in the problem, then use a…
Solution: There are 8 (= 2 * 2 * 2) different cases, and the results are provided in the following table: | n | σ | 1-α | 1-α/2 | Zα/2 | σZα/2 √n | Confidence Interval | |-----|-----|-----|-------|------|----------|----------------------| | 10 | 0.01| 95% | 97.5% | 1.96 | 0.062 | (Xn…
Question 3 10pts A test was given to a group of students. The grades and gender are summarized in the chart below. If one student is chosen at random from those who took the test, find the probability that the student was male given they got a B. Write your answer as a decimal rounded to 4…
A student-faculty government committee of 4 people is to be formed from 21 student volunteers and 4 faculty volunteers. Find the probability that the committee will consist of the following, assuming the selection is made at random: A student-faculty government committee of 4 people is to be…
accompanying table. able[[,y(x,y),0,5,10,15],[,0,0.01,0.06,0.02,0.10,,,,,],[x,5,0.04,0.13,0.20,0.10,,,,,],[,10,0.01,0.15,0.17,0.01,,,,,]] (a) Compute the covariance for x and Y. (Round your answer to two decimal places.) Cov(x,Y)= (b) Compute ho for x and Y. (Round your answer to two decimal…
Suppose the manufacturer can employ a temporary employee to work more hours in step S2 at $10 for each additional hour. Is it worthwhile? How much would the total cost change for one more hour? Group of answer choices It is worthwhile as it could save $7.5 for each additional hour. It is not…
1024,256,64,16,4, (Type an integer or a simplified fraction.) Determine the most probable next term in the list of numbers 1024,256,64,16,4 1024256,64,16,4, (Type an integer or a simplified fraction.)
Test whether β(0) is significant at α = 0.05. The p-value is 0.05. H0: We conclude that the intercept of the regression significantly contributes to the annual maintenance cost. What is the interpretation of β(0)? Is the interpretation of β(0) reasonable? The interpretation is because β(0) has…
Click the icon to view the table. Percentage of scores GREATER than 122 is ◻%. (Round to two decimal places as needed.) Data table Standard Scores and Percentiles for a Normal Distribution Full data set able[[Standard score,%, able[[Standard],[score]],% Use the normal distribution of IQ…
c. mu_{(D)}>0 d. mu_{(D)<0 b. What is the alternative hypothesis? a. mu_{(D)}=0 b. mu_{(D)} eq0 c. mu_{(D)}>0 d. mu_{(D)}<0 Data: able[[Action Movie,Horror Movie],[11,12],[15,18],[16,17],[20,21],[24,22],[25,23],[26,29],[13,16],[18,21],[19,23],[20,15],[24,25]] c. What are your degrees of…
Based on your sample, you will conduct a hypothesis test to test two of the claims of the above article. Using the same Excel sheet as last week, answer the following in the "Week 6" tab: Claim: The average age of online students is 32 years old. Can you prove it is not? - What is the null…
Hi, I will help you with this. Let's start by correcting the errors in the text provided: Original: (a) Show that P(x<=n)=1-(1-p)^(n),ninZ^(+). (b) Deduce an expression for p=0.2n The random variable X has a geometric distribution with parameter p (a) Show that P(Xn)=1-(l-p)", neZ*. [3 marks]…
A small-business owner collected data from a random sample on the amount spent on an order and the time it takes to fill the order. The resulting data were used to complete a linear regression analysis of time, in minutes, versus amount spent, in dollars, and to create the following residual…
Dr. Rankin was interested in examining the impact of mindfulness meditation on reducing stress levels. After collecting data, she calculated the mean and standard deviation for reduction in stress scores among participants. Further, she wanted to compare the collected data to a known population…
Can you also show a pixture od the answrs in excel Compute the correlation between motivation and GPA. Test for the significance of the correlation coefficient at the .05 level using a two-tailed test. True or false? The more highly you are motivated, the more you will study. Which did you…
Question 2: ***DISCRETE MATH*** You are celebrating the end of the semester at your local pub. Being an intrepid computer scientist, you run a randomized algorithm to determine what you order. Every time the wait staff asks you your order, you roll 2 fair, six-sided dice and add them together…
A machine that fills beverage cans is to be replaced by a new system if the variance in the new machine has a variance of less than 0.01 when filling up 12 ounces of beverage in a can. In order to test the variance in the new machine 10 cans are filled and the amount of liquid in each can is…
Calculate the minimum building airflow for the well-shielded, two-story residence. Building data: Living space area = 2400 ft^2 Number of occupants = 5 Number of stories above grade = 2 Location in Houston, TX egin{table}[h] centering egin{tabular}{|c|c|c|c|c|c|} hline Zone & Number of…
This data is from a sample. Calculate the mean, range, standard deviation, and coefficient of variation. Round answers to two decimal places. able[[x This data is from a sample. Calculate the mean, range, standard deviation, and coefficient of variation. Round answers to two decimal places. X…
Assume the random variable x is normally distributed with mean μ = 50 and standard deviation σ = 7. Compute the probability. Be sure to draw a normal curve with the area corresponding to the probability shaded. OA. P(34 < X < 57) probability shaded. 0 Which of the following normal curves…
(a) The transition probability of a Markov chain with threc states {0,1,2} is P=([(3)/(4),(1)/(4),0],[(1)/(4),(1)/(2),(1)/(4)],[0,(3)/(4),(1)/(4)]) and the initial state distribution of the chain is P{x_(0)=i}=(1)/(3),i=0,1,2. Find (i) P{x_(2)=2} and (ii)…
Show all steps with answers please. QUESTION 22 For the following data: Group 1 Group 2 8 3 5 4 6 9 4 8 2 7 Calculate the following: a.Ss 1 b.SS 2 c.SS D d. e.s
A journal on mental health presented the results of a survey of over 6,000 clients of a state's mental health department. One of the many variables measured for each mental health patient was frequency of social interaction (on a 5-point scale, where 1 = very infrequently, 3 = occasionally, and…
a. Using the results in column (1), what is the expected change in price of building a 500-square-foot addition to a house? Construct a 95% confidence interval for the percentage change in price. b. Comparing columns (2) and (3), is it better to use Size or ln(Size) to explain house prices?
Chapter 8: (8.2) Confidence Intervals (Mean) - sigma unknown - Homework Score: 13.8/16 9/10 answered Progress saved sqrt(0)◯,vdots (0)/(2) pts Details Assume that a sample is used to estimate a population mean mu . Find the 98% confidence interval for a sample of size 35 with a mean of 26.8…
You wish to test the following claim (HA) at a significance level of α = 0.10. Ho: P1 = P2 HA: P1 > P2 You obtain 506 successes in a sample of size n1 = 720 from the first population. You obtain 361 successes in a sample of size n2 = 543 from the second population. What is the test statistic…
Use the Standard Normal Table or technology to find the z-score that corresponds to the cumulative area 0.0071 . z= (Type an integer or decimal rounded to two decimal places as needed.) Use the Standard Normal Table or technology to find the z-score that corresponds to the cumulative area…
[20 points] Consider the problem of predicting how well a student does in her second year of college/university, given how well she did in her first year. Specifically, let x be equal to the number of "A" grades (including A-. A and A+ grades) that a student receives in their first year of…
It is known that the number of misprints on a textbook page follows a Poisson distribution with a mean of 0.004 . Which of the following is the standard deviation of the number of misprints on a textbook of 475 pages? A. 0.004 B. 0.063 C. 1.9 (e^(-lambda )*lambda ^(x))/(x!) D. 1.378 X! A.…
Which of the following come out to about the same number when they are calculated with the same set of data? a. mean absolute deviation and variance b. mean absolute deviation and standard deviation c. variance and standard deviation 10. Which of the following come out to about the same number…
Exercise 7. Let xi be a real random variable whose c.d.f is F_(xi ). Find the c.d.f's of the following random variables in terms of F_(xi ) : zeta =|xi |. eta =ln(xi ) (assuming xi >0 ). heta =exp(xi ). 1 Suppose that xi is normally distributed with mean mu and variance sigma ^(2), i.e its…
A psychologist wanted to know whether there were visual spatial differences in the development of children if they are raised within a hunter-gather culture. A sample of n=36 children between the ages of 6 - 8 years old were recruited and a cognitive test scores 145 . The general mu =120 with a…
Question 2 (2 points) You'd like to be able to compare scores on the SMART with another standard aptitude test -- the ACT. In order to do this, we need to transform the SMART distribution so that it has the same mean and standard deviation as the ACT. ACT has a mean of 20 and standard deviation…
You are playing a game with a friend. You roll a die. If you roll a 3 , you win $15. If you lose, she collects $3. Is this a fair game? Explain. 7. You are playing a game with a friend. You roll a die. If you roll a 3, you win $15. If you lose, she collects $3. Is this a fair game?Explain.
Calculate the Sensitivity of a screening test having these results: Of those with the disease: 1000 test positive, 100 test negative; of those without the disease 250 test positive and 2500 test negative. Answer as a percentage. Calculate the Sensitivity of a screening test having these…
An SRS of 50 students were selected from a population of 200 students. The population has a parameter equal to 0.85. What is the center and spread of the sampling distribution? In this example when calculating spread of sampling distribution do you divide by sample or by…
Pickmeup is a drive-through coffee house that has room for 3 cars in the driveway. The line cannot exceed 3 cars, and because they are exclusively drive-through, customers may be turned away. In the morning, the arrival rate is 40 cars per hour (Poisson distributed) and with the servers working…
Sample standard deviation S2 = 6. In performing the pooled-variance t-test, how many degrees of freedom are there? There are __ degrees of freedom. (Simplify your answer.) Assume there is a sample of n = 5, with the sample mean X1 = 35 and a sample standard deviation of S1 = 3, and there is an…
(1 point) Match the correlation coefficients with their scatterplots. Select the letter of the scatterplot below which corresponds to the correlation coefficient. (Click on image for a larger view.) r=0.22 r=0.89 r=-0.74 r=-0.49 4.1 Scatter Diagrams Linear Correlation: Problem 5 Previous…
According to a Pew Research Center project, the main sources of news information for Americans are television and:
Provided below are summary statistics for independent simple random samples from two populations. Use the pooled t-test and the pooled t-interval procedure to conduct the required hypothesis test and obtain the specified confidence interval. ar{x} _(1)=17,s_(1)=6,n_(1)=19,ar{x}…
Part 6 of 9 Points: 0 of 1 The regression model hat(Y)_(i) = -349.3 + 1.7513x_(1i) - 0.0912x_(2i) predicts standby hours based on total staff present, x_(1i), and remote hours, x_(2i), for week i. The data from which the model was constructed are provided. Use the regression model and the data…
What would be quantitative analysis be for the mathematical results of warning and actions Production 000000 Week1 Week2 Estimated Lost Production Week3 Week4 Mean Weeks Week5 UCL Week 6 LCL Week7 Week8 Control Chart LCL Mean UCL
Compute the probability of x successes, using the binomial formula. a. n=6,x=3,p=0.03 b. n=4,x=2,p=0.18 c. n=5,x=3,p=0.63 Compute the probability of X successes,using the binomial formula. a.n=6,X=3,p=0.03 b.n=4,X=2,p=0.18 c.n=5,X=3p=0.63
Age discrimination: The following table presents the numbers of employees, by age group, who were promoted, or not promoted, in a sample drawn from a certain industry during the past year. [ egin{array}{|c|c|c|c|c|} hline ext{Age} & ext{} & ext{Under 30} & ext{30-39} & ext{40-49} &…
In a board game, a player rolls a 6 sided dice with the numbers 1 to 6 on it, and spins a spinner with the letters A to D on it. a. Create an outcome table to list all the possible outcomes. (1 mark) able[[,1,2,3,4,5,6],[A,,,,,,],[B,,,,,,],[C,,,,,,],[D,,,,,,]] N a.s with the letters A to D on…
2. You believe that the mean number of cats present in Guney Campus Manzara at any given moment is 4. Please answer each part below separately. a. You have counted the number of cats 9 times, and ar{x} turned out to be 3.3. If a frequent visitor of the place told you that sigma^2 for the…
You are a financial analyst who is evaluating the quarterly sales of two different companies. Based on the following data (numbers are in millions of dollars), construct an ANOVA table, state your null and alternate hypotheses, and indicate your conclusion based on a 0.05 level of…
Which of the following statements best describes current conservative attitudes about immigration? a) There is no division of opinion among conservatives because almost all conservatives strongly oppose immigration. b) There is no division of opinion among conservatives because almost all…
The data below are the temperatures on randomly chosen days during a summer class and the number of absences on those days. Calculate the correlation coefficient, r. able[[Temperatur, x,75,88,94,93,91,101,78,103,83],[Number of absences, y,5,9,12,12,10,17,6,17,7]] A. 0.819 B. 0.890 C. 0.980 D.…
Data is collected to determine if the proportion of occupants in cars who wear seat belts (p_(1)) is less than the proportion of occupants in pickup trucks who wear seat belts (p_(2)). Which hypotheses are appropriate to test…
b. What is the probability that weekly demand will be greater than the number of available radios? The probability is 0.10 c. What is the expected weekly profit from the sale of the alarm clock radio? (Remember: There are only five clock radios available) The expected weekly profit is $◻.…
Calculate A_(x) in terms of u,y, and z. Problem 4 Given: (10 points) (i) Ax:n= u, (iii) Ax+n = z. Calculate Ar in terms of u, u, and z
For arbitrary random variables x and Y, prove that Cov(x,E[Y|x])=Cov(x,Y). 4.For arbitrary random variables X and Y,prove that Cov(X,E[Y|X])=Cov(X,Y)
QUESTION #9 (15 marks) How can engaged couples improve their chances of a happy marriage? A study selected 19 couples and tested the quality of their communication skills 3 months before marriage. Leading up to the marriage, the couples were then asked to participate in a communication skills…
*DISCRETE MATH* Question 4: We roll a fair red die and a fair blue die. We define the following random variables. R={(1 if the red die is even ):} 0 otherwise B={(1 if the blue die is 3 or less ):} 0 otherwise S={(1 if the sum of both dice is even ):} 0 otherwise It should be clear that R…
A two-echelon distribution network [one central distribution center (x) and two stores A and B)] has 120 units of a product at the central distribution center, and no additional receipts are expected at the central distribution center for at least a week. The on-hand inventory and…
10. Events A and B are such that P(A) = , P(B) = and P(A|B) = 11. (a) Find P(A ∩ B). (b) Find P(A ∪ B). (c) State with a reason whether or not events A and B are independent.
Exercise 10-11 CAS Exam 3 Spring 2005 #22 Drivers are classified as "good" or "bad." Results of the classification are assumed to be binomially distributed, with probability of being "good" equal to p. A sample consists of 100 drivers. Determine the critical value for testing the hypothesis…
Question 3 [1.0 pt: a (0.4 pts), b (0.6 pts)] Imagine we want to estimate the population mean of some variable Y (e.g. earnings). To keep things simple, think of taking a random sample of size two from this population. Write Y_(1) and Y_(2) for the two values you get from the sample. Here are…
Example 10-6 A quality inspector counted the number of defective parts a prototype machine made in samples of 100 taken every hour during a 20-hour period. a) Using the following data, construct a Three Sigma control chart for the sample proportion of defectives. b) Is the machine producing a…
Determine the area under the standard normal curve that lies between (a) Z=-1.35 and Z=1.35, (b) Z=-0.77 and Z=0, and (c) Z=-1.65 and Z=1.18. (a) The area that lies between Z=-1.35 and Z=1.35 is (Round to four decimal places as needed.) Round to four decimal places as needed.
A researcher conducted a left-tailed hypothesis test on the amount of caffeinated coffee drunk by coffee drinkers, with HA: μ < 23. They found the p-value to be 0.0799, using a significance level of α = 0.10. Select the 2 best answers: Select all that apply: - Fail to reject the Null…
Data: 198,188,208,205,210,198,189,183,190,195,176, 189,174,155,180 c. What are your degrees of freedom? d. What is the critical t value ( t table)? e. What is the obtained t value? f. What can you conclude? a. Reject the null hypothesis. American men at a southwestern college are significantly…
A.) Show that k=2/9 B.) Find the mean and standard deviation of x
Bill keeps the phone numbers for his seven closest friends(three men and four women) in his digital phone memory. In how many ways can he list them if (a) men are listed before women? (b) men are all listed together? (c) no two men are listed next to each other? 7. Bill keeps the phone numbers…
(5 points) The five-number summary for a set of data is as follows: Minimum: 25, Q1: 32, Median: 40, Q3: 47, Maximum: 60. If there are 250 values in the data set, about how many values are between 32 and 47? (Hint: Think about percentiles.) 5.5 pointsThe five-number summary for a set of data…
Fifty people participated in a medical study. Each participant contributed two eye measurements (labeled y ). Thus the study had a total of 100 measurements on y from the 50 participants. It is suspected that the two eye measurements from the same person are correlated. Without sufficient…
Confidence Intervals for One Proportion Worksheet Math 181 Statistics Helgeson Due at the start of class on Monday, April 1, 2024. On the Sections 2.1, 2.2 & 2.3 (Theory-Based Hypothesis Testing for a Single Variable) lecture notes, we used simulation and also a theory-based hypothesis test…
6 and 7) Question 2 of 15 This question: 1 point(s) possible Survey results indicate that σ = 15.8 books. A 99% confidence level requires subjects. (Round up to the nearest whole number as needed.) 6 and 7) Question 2 of 15 This question: 1 point(s) possible People were polled on how many books…
In a class with 40 students, 18 are boys, and 22 are girls. In the last exam, 2 boys and 2 girls achieved grade A. Calculate the probability of, in a random selection of two people, a boy or a student with grade A being selected. (3)/(4) (1)/(2) (4)/(40) (18)/(40) In a class with 40 students,…
Coefficient to assess the normality of the data. Click here to view the data table. Click here to view the table of critical values for normal probability plots. Data Table (b) The correlation is (Round to three decimal places as needed.) able[[Index,…
What is the probability of randomly selecting a respondent who is more conservative than "moderate"? What is the probability of randomly selecting a respondent who is either "slightly liberal" or "slightly conservative"? What is the probability of randomly selecting two respondents, the first…
Question 8 (1 point) Assume we want to run a hypothesis test for the population mean. If we were to find a p-value of 0.0431 , what decision should we make at the 0.05 level of significance? Don't reject the null. There is NOT evidence of the alternate. Accept the null. There is evidence of the…
Two dice are rolled, find the probability of on which the sum of the dice is less than 18. Two dice are rolled,find the probability of on which the sum of the dice is less than 18
To understand the numbers in this article, you need to understand "median wage". To figure out the median wage, you would list the wages in order from smallest to largest and then find the wage where half of the people make less and half of the people make more (50th percentile). This is the…
A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 237.1-cm and a standard deviation of 2.2-cm. Find the probability that the length of a randomly selected steel rod is greater than 240-cm. P(x>240-cm)= Enter your answer as a number accurate to…
Use the following data set (data file provided) to determine if a significant statistical difference between 3 groups in the following experiment. Please perform a One-Way ANOVA analysis (ANOVA Single Factor), and then interpret the data. In this experiment, subjects perform a 1-hour…
. From the statistical and stochastic test scores of 100 students, 40 samples were randomly selected, namely: able[[,53*,57*,62*,65,71*,74,83],[42*,54*,58,62,65,72*,75,86],[45*,55,58,64,68*,73,76,86],[49*,56*,58,64,70*,73,78,86],[50*,56*,60,64*,71*,74,82,87]] a. What is the lowest value that…
Bag A contains 2 red and 3 green balls. a) Two balls are chosen at random from the bag without replacement. Find the probability that 2 red balls are chosen. Bag B contains 4 red and n green balls b) Two balls are chosen without replacement from this bag. If the probability that two red balls…
Hi need Help ASAP, thank you! Anil is a candidate in an election. He received 40% of the votes. A random sample of 120 voters is chosen. Use an approximation to find the probability that, of the 120 voters, between 36 and 54 inclusive voted for Anil. Anil is a candidate in an election. He…
For the following information, determine whether a normal sampling distribution can be used, where p is the population proportion, alpha is the level of significance, hat(p) is the sample proportion, and n is the sample size. If it can be used, test the claim. Claim: p>=0.34;alpha =0.04.…
A politician claims that the mean salary for managers in his state is more than the national mean, $82,000. Assume the population is normally distributed and the population standard deviation is $7600. The salaries (in dollars) for a random sample of 30 managers in the state are listed. At α =…
A population of values has a normal distribution with mu = 240 and sigma = 71.8. You intend to draw a random sample of size n = 42. Find the probability that a single randomly selected value is less than 217.8. P(X < 217.8) = Find the probability that a sample of size n = 42 is randomly…
QUESTION 5 What happens to the length of the confidence interval as the standard error increases? The length of the confidence interval increases. The length of the confidence interval decreases. There is no change to the length of the confidence interval as long as alpha doesn't…
Given two independent samples, conduct a hypothesis test for the difference between the means. Consider a significance level of 0.05 . Use a confidence interval to decide whether to reject or not H_(0). H_(0):mu _(1)-mu _(2)=0 H_(1):mu _(1)-mu _(2)!=0 able[[Sample (1),Sample…
Question 4 Use the data in the following table, which lists survey results from undergraduate school students. Assume that subjects are randomly selected from those included in the table. If one of the students is randomly selected, find the probability that the selected student took online…
Ignore the questions that talk about RStudio (the questions with yellow highlighter) In the class of 2021, where scores for all students who took the SAT during high school. For Problem Set 6, you have 524 students. 1. What is the probability that a student has a SAT Math score lower than…
Fill out the combined proxy column 7.7 + (3.7 x Production Cost) Fill out the combined proxy column 7.7 + (3.7 x Production Cost) + (7.6 x Promotion Costs) + (0.8 x Book Sales) Box Office Receipts Production Costs Promotion Costs Book Sales Combined Proxy 85.1 8.5 5.1 4.7 81.6 106.3 12.9 5.8…
A survey collected data about the number of TVs that a family of four has in their household. The number of TVs is distributed as shown. Find the mean, variance, and standard deviation for the distribution. (Write answers rounded to one decimal place. Use the letter A or B to answer the…
Find the EZ and Var(Z) for the Z∼(mu ,sigma ). Find the EZ and Var(Z) for the Z~(u,o)
The hypothesis is that there is no difference in the proportion of all men and women in the city who intended to vote for Trump. The null hypothesis is that the proportion of men intending to vote for Trump is equal to the proportion of women intending to vote for Trump. The alternative…
computers as factory C.The probability that a computer produced by factory A is defective is 0.04, the probability that a computer produced by factory B is defective is 0.033, and the probability that a computer produced by factory C is defective is 0.046. What is the probability it came from…
What is the correct sequence of events (oldest to youngest) for this diagram? Enter the letters with no spaces or commas. For example: ABCD What is the correct sequence of events (oldest to youngest) for this diagram? Enter the letters with no spaces or commas. For example: ABCD B c D A I
Problem 1 If x_(1),...,x_(n) are iid from Unif(- heta , heta ) RVs. Find a consistent estimator of heta . Problem 1 If X.....Xn are iid from Unif(-0,9) RVs. Find a consistent estimator of 0.
The analysis of results from a leaf transmutation experiment (turning a leaf into a petal) is summarized in the table by type of transformation completed. Let A denote the event that a leaf completes the color transformation, and Let B denote the event that a leaf doesn't complete the textural…
Model 1 (A) Use a simple linear regression of y= Price against x= Odometer. State your NULL & ALTERNATE hypotheses (for each of these models) and your conclusions. Use α = 0.05 throughout this project (for each of these models). Plot the graph of price against odometer and add it to your…
Consider a game where 3 contestants each roll an n-sided die. On the first play of the game, each will roll a 3-sided die. On the next play of the game, each contestant will roll a 4-sided die. On each subsequent play, the number of sides on the die will increase by 1. a) Suppose on any play…
It's New Year's Day, 2023, and thanks to a long night out with friends, you are off to a slow start. Little Cottonwood Canyon is closed due to extreme avalanche conditions. You decide to check out the weather at Snowbasin. Snowbasin's Boardwalk observing site provides snow interval observations…
ch babies first crawl. [ egin{array}{|c|c|c|c|c|c|c|} hline 49 & 31 & 46 & 34 & 39 & 27 \ hline 49 & 39 & 54 & 26 & 37 & 28 \ hline end{array} ] Chi-Square Distribution Critical Values Table Click the icon to view the table of critical values of the chi-square distribution Select the correct…
SUMMARY OF FORMULAS FORMULA 6.1 FORMULA 6.2 Confidence interval for a sample mean (large sample, population standard deviation known): Confidence interval for a sample mean (large sample, population standard deviation unknown): Confidence interval for a sample proportion (large…
For a standard normal distribution, find: P(0.39 < z < 2.18)
The size of fish is important to commercial fishing. A study conducted in 2102 found that the length of Atlantic cod caught in nets in Karlskrona, Sweden is normally distributed with mean 49cm and standard deviation 4cm. Source: Condition indices of Atlantic cod (Gadus morhua) biased by…
5.2 Some Probability Rules Compound Events: Problem 16 Results for this submission At least one of the answers above is NOT correct. ( 1 point) In a certain community, 35% of the families own a dog, and 20% of the families that own a dog also own a cat. It is also known that 29% of all the…
Let {Y_(t)} be a stationary process with model equation Y_(t)=24+(2)/(3)Y_(t-1)+epsi lon_(t)+(1)/(2)epsi lon_(t-1),tinZ, where {epsi lon_(t)} is a white noise process such that the epsi lon_(t) are NID(0,sigma ^(2)). Let y_(1),y_(2),dots,y_(T) denote an observed realization of the process up to…
Suppose we estimate the regression y = β0 + β1x1 + β2x2 + u, using 282 observations. We then conduct several hypothesis tests with a significance level α = 0.05: F-test for goodness of fit, two Breusch-Pagan tests, and two White tests (using the predicted dependent variable in a cubic…
PRACTICE ANOTH The method of tree-ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution. able[[1194,1292,1292,1250,1268,1316,1275,1317,1275]] (a) Use a calculator with mean and standard…
Suppose X1, . . . , Xn are iid RVs with mean lambda and variance lambda . Find a large sample 95% confidence interval for √lambda .Problem 2 If x_(1),dots,x_(n) are iid Unif(0, heta ). Find a large sample 95% confi- dence interval for heta . Problem 2 If X1,..:,Xn are iid Unif(0,0). Find a…
The Normal Approximation to the Binomial Distribution Consider the probability that less than 14 out of 157 flights will be delayed. Assume the probability that a given flight will be delayed is 12%. Approximate the probability using the normal distribution. Round your answer to three decimal…
2 Question 2 able[[Age,17,21,64,3 r.,86,41,26,58,46],[Income,10,500,55,367,45,678,37,985,65,786,126,121,97,360,67,587,278,941]] Sean is a 27 year old making an income of $85,500. Is he making above the predicted value of 2 Question 2 Age 17 21 64 86 41 26 58 46 Income 10,500 55,367 45,678…
The sample mean and standard deviation from a random sample of 29 observations from a normal population were computed as ar{x} = 28 and s = 15. Calculate the t statistic of the test required to determine whether there is enough evidence to infer at the 4% significance level that the population…
Interviewed in a random order. (a) What is the probability that x of the top five candidates are interviewed on the first day? h(x; 7, 5, 11) h(N; 4, 5, 11) h(N; 7, 11, 5) h(x; 4, 11, 5) (b) How many of the top five candidates can be expected to be interviewed on the first day? (Round your…
Assume that a person spins the pointer and is awarded the amount indicated by the pointer. Determine the person's expectation. A. $4.50 B. $3.25 C. $6.50 D. $4.25 To determine the person's expectation awarded the amount indicated by the pointer, assume that a person spins the pointer and is…
Question 6 7 pts According to admissions data from the Massachusetts Institute of Technology (MIT), 4.1% of freshman applicants are admitted to the college. If 33,240 individuals applied to MIT last year, find the probability that at least 1,400 of them were admitted. Round your answer to 4…
We have this population: {2,3,4,5,6}. The mean is 4 , but it is ignored by the investigator. He still insists to estimate the mu from samples. We have three possible estimators which are: Estimator 1: sample mean Estimator 2: sample median Estimator 3: average of the largest and the smallest…
Problem 3 Suppose x_(1),dots,x_(n) are iid RVs with mean lambda and variance lambda . Find a large sample 95% confidence interval for sqrt(lambda ). Problem 3 Suppose X..... Xn are iid RVs with mean X and variance A. Find a large sample 95% confidence interval for VX
Two men A and B toss a coin in turn until a head occurs. The man who tosses a head wins. a) Find the probability that A wins. b) If the coin is biased and head occurs with probability p, what is the probability that B wins? c) Let X be the number of trials required to produce the 1st success,…
In Box I are bulbs that are guaranteed to last one year, selling for R3 each. In Box II are bulbs with no guarantee selling for R1 each. The store owner explains that Box II came to him unlabeled. He knows that all of the bulbs in Box II are of one type and declares that it is equally likely…
A single sample of 9 data points is taken. If you perform a right-tailed one sample hypothesis test for the means with an alpha value of 0.05, what is your degree of freedom value and what is the t-critical value for this test.
The correct answer is: c. X2 < 29.1413
When it comes to people snoring while sleeping, many will use an anti-snoring device. About 70% of users lose 10 pounds in two weeks. Create a hypothesis test to evaluate this claim. Represent it as a null hypothesis and an alternative hypothesis.
Let G be a set containing 5 elements a, b, c, d, and e. Under a binary operation called the multiplication, we know that ab = c and cd = dc = e. Show that G is not a group under this operation
Find the critical values chi Subscript Upper R Superscript 2 and chi Subscript Upper L Superscript 2 for the given confidence level c and sample size n. c equals 0.80, n equals 16 Question content area bottom Part 1 chi Subscript Upper R Superscript 2 equals    enter your response here (Round…
9. (Normally distributed quality characteristic) Suppose a quality characteristic of a product is normally distributed with mean 18 and standard deviation 1.5 . The specification limits furnished by the customer are (15,21). Determine what percentage of the product meets the specifications…
Find the regression line for data points X | Y 85 |80 90|40 100| 30 102| 26 105| 20
Consider a continuous random variable X with probability density function f(x) = c · x(1 − x)^5, 0 ≤ x ≤ 1,
A random sample of 12 cats was collected to study the lifespan with the data listed below. Find the 87% confidence interval for the population mean. Round all values to two decimals. 13, 15, 18, 16, 8, 12, 13, 15, 8, 13, 12, 11. Provide your answer below.
a) A simple random sample of size 20 has mean x = 3.19. The population standard deviation is a = 1.51. The population is approximately normally distributed. Construct a 90% confidence interval for the population mean.
What is the empirical probablity of obtaining 9 heads in 10 tosses
Project for MA235 STEP 1: You are doing research on a topic of interest. Decide what you want to research. To complete this research you will need to determine a question that has a numerical response. Identify a random sample of n = 15 men and n = 15 women. NOTE: Though it is suggested, you do…
I need some help interpreting some data. I have some of the answers but I am having a hard time finishing this. 1. Describe the method used and why you used the method. We will apply this by sending out emails of our survey to 300 random students at BYU-Idaho. We chose this method because we…
Ciorstaidh's investment has a 20% chance of producing a 14% return, a 50% chance of producing a 9% return, and a 30% chance of producing a -2% return. What is its standard deviation? a. 4.46% b. 5.81% c. 6.00% d. 0.36%
Find the regresion line: X: 40, 60, 80, 100, 120, 140, 160. Y:0, 0.2, 0.4, 0.6, 0.8, 1, 1.2
Let be a random variable with a standard normal distribution. Find the indicated probabilities. a) P (z<-1.32)= ____________ c) To the left of z=72, _____________________ b) P(z<=-2.15)=__________________ d) P(z>=2.17) =_____________ e) P(z>=-1.50)= ___________
Describe a forensically relevant scenario that might involve a binomial random variable and thus require the use of the Binomial distribution
The probability distribution that can be described by just one parameter is the _____ distribution. a. uniform b. continuous c. normal d. exponential
If lambda is set at 34, what is the probability of 29 or more but no more than 46?
a) Suppose there is population with 4 customers. The interested categorical variable is the proportion of customers that prefer your brand. In this population, one of them prefers your brand, and the others do not. (i) Find the population proportion (pi ) if customers that prefer your brand. …
I need the R codes and the answers to the question. Using R Studio, load wage1 data from the wooldridge package. Estimate the model log(wage) = B0 + B1educ + B2exper + B3tenure + u. Interpret the coefficients and the model overall. Calculate the VIF values. Do you think the multicollinearity is…
As a part of the study of the development of thymus glands, researchers weighed the glands of chick embryos after 14 days of incubation. The thymus weights (mg) were as follows: 29.6, 30.1, 24.7, 21.5, 33.4, 29.8, 41.5, 32.9, 33.7, 28.0, 34.6, 44.9, 37.8, 23.4, 27.1, 41.5, 43.2 a) Calculate…
According to a survey, the mean student loan at graduation is $50,000. Suppose that student loans are normally distributed with a standard deviation of $5,000. A graduate with a student loan is selected at random. Find the following probabilities. (Round your answers to four decimal…
Here's a list of units of toys sold by T&K over a 12-month period during 2015. Were the sales of 31,456 for one month in 2016 significantly different from monthly sales in 2015? Units Sold in 2015 January 34,518 February 29,540 March 34,889 April 26,764 May 31,429 June 29,962 July …
A bag contains 3 boxes of sour candy and 5 boxes of chocolates. (a) If you draw one box at random, what is the probability that it is a box of sour candy? (b) A friend asks for a box of chocolates and removes it from the bag. What is the probability now of drawing a box of sour candy at random?
Using the interquartile range which of the following are outliers select all correct answers 43 4448 6470
A random sample of 5 patients has an average time for a kidney stone pass of 10.2 days and a sample standard deviation of 0.39. Find the 89% confidence interval for the population mean. Round all values to two decimals. Provide your answer below.
Why is an understanding of different scales of measurement(nominal, ordinal, interval and ration) important for the purpose of questionnaire design?
A random sample of six nursing students has an average amount of sleep in 24 hours of 7.6 hours and a sample standard deviation of 3.78. Find the 79% confidence interval for the population mean. Round all values to two decimals. Provide your answers below.
7 Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 5.8-in and a standard deviation of 0.9-in. In what range would you…
Part 3: Hypothesis Test: (2 points) I had recently conducted an experiment: the moment the fuel light in my car goes on, I will find a closest gas station to fill up the tank. But the number of gallons needed to fill up are all over the place. This makes me wonder how we can verify it is…
A firm specializing in agricultural products wants to conduct a test-market offering for one of its new products. A random sample of 800 potential customers is chosen to receive advertising material describing the new product, and it was determined that 600 out of the 800 people made a…
If, in a sample of n = 30 selected from a normal population, X̄ = 52 and S = 12, what are the critical values of t if the level of significance, α = 0.05, the null hypothesis, H0, is μ = 50, and the alternative hypothesis, H1, is μ ≠50?
A distribution of values is normal with a mean of 124.5 and a standard deviation of 4.7. Find the probability that a randomly selected value is greater than 138.1. P(X > 138.1) =
What is the probability that a randomly selected baby born this year weighed 3000 to 3999 grams or was postterm
For a standard normal distribution, find: P(z > -2.72) Express the probability as a decimal rounded to 4 decimal places.
In a bag there are 10 balls. There are 4 red balls and the remaining balls are green balls removed at random and the color noted, the ball is replaced. Then a second ball is removed at random and the color is noted. P(1 red ball and 1 green ball picked)= 4/10*6/10=6/25 P(2 red balls picked…
5.40. A multiple-choice test consits of eight questions and three answers to each question (of which only one is correct). If a student answers each question by rolling a balanced die and checking the first answer if he gets a 1 or 2, the second answer if he gets a 3 or 4, and the third answer…
If two events A and B are such that P(A) = 2/3, P(B|A) = 1/2 and P(A') = . Use this information to answer questions 5-11. 5. Find P(A|B'). 6. Calculate P(A|B). 7. Calculate P(A'|B). 8. Calculate P(A'|B'). 9. Find P(A ∩ B'). 10. Find P(A' ∩ B). 11. Find P(A' ∩ B'). 12. Suppose that A and B are…
An agent sells life insurance policies to 6 equally aged, healthy people. According to recent data, the probability of a person living in these conditions for 25 years or more is 2/5. Calculate the probability that after 25 years: A.) Find the probability that all 6 people are still living. (10…
Tree heights: Cherry trees in a certain orchard have heights that are normally distributed with mean mu =117 inches and standard deviation sigma =1 4 What is the probability that a randomly chosen tree is between 90 and 102 inches tall?
Read the following question about survey data collection and provide a brief response Bruce, a biology professor, is interested in opinions about the summer research program at his college and decides to conduct a survey of students who've participated. He makes the survey available online and…
A researcher conducted a study with 12 patients and found the sample mean of the kidney stone passing to be 6.4 days with a sample standard deviation of 3.4. They want to test if this is greater than the regional average of 9 days at a significance level of a=0.25. What is the p-value for this…
Select the 99% CI of mean differences using the following: df = 9 SE = 4.2 Mean difference = 5.5 a. (-10.16, 15.16) b. (2.21, 6.31) c. (-8.15, 19.15) d. (-7.22, 18.14)
Find the area under the standard normal curve that lies between z = -1.28 and z = -0.29.
suppose over the course of one hour, a player can expect to be dealt about 40 hands. how much should a player expect to win or lose over the course of three hours
For the following questions, do not compute numerically. (a) Compute the number arrangements (i.e. permutations) of 52 cards in a standard deck, in which red cards (hearts and diamonds, not necessarily in that order) appear before black cards (spades and clubs, not necessarily in that…
A random sample of 420 observations produced a sample proportion equal to 0.32. Find the critical and observed values of z for the following test of hypotheses using α = 0.1. Ho: p = 0.30 versus H1: p ≠0.30. Round your answers to two decimal places. Zcritical left = Zcritical right = Zobserved
A) For a binomial distribution with a sample size equal to 4 and a probability of a success equal to 0.18, what is the probability that the sample will contain exactly two successes? Use the binomial formula to determine the probability. Question content area bottom Part 1 The probability that…
d¯ = -1.000 sd = 4.899
A driver's reaction time to visual stimulus, R, is normally distributed with a mean of 0.4 seconds and a standard deviation of 0.05 seconds.a. Create the probability density function of R.b. Determine the mean, variance, and standard deviation of R.c. Plot the probability density function of…
A university has admitted a group of n students with varying skill levels. To better accommodate the students, the university has decided to create classes tailored to the skill levels. A placement examination will return a skill level that will be used to group the students, where levels[i]…
The following products are included in the current State Farm Health product portfolio EXCEPT: -Short Term Disability Insurance - Long Term Disability Insurance -Long Term Care Insurance -Medicare Supplement Insurance The monthly benefit range for the Short Term Disability Insurance policy…
Students are given a multiple choice exam with 4 questions. The random variable X = the number answers correct. Possible values = {0, 1, 2, 3, 4}. From past data, 10% of students get zero correct answers, 10% get exactly one correct answer, 20% get two correct, and 40% get three correct. It can…
Problem 4. (15 points) Let X denote the lifetime of a radio, in years, manufactured by a certain company. Assume that the lifetime is exponentially distributed, with a mean of 5 years. (a) (5 points) What is the probability p that a radio lasts more than 15 years? (b) (5 points) What is the…
Customers arrive at a pharmacy to be served according to a Poisson process with rate lambda . Knowing that q <= n, m <= n, and q + m <= n, and that 0 < t1 < t2 < t: (a) (3 points) If n customers arrive in the interval [0, t], what is the probability that exactly q customers have arrived in the…
In a laboratory where the temperature is being kept at the range of 20 degrees and 25 degrees it is found that over the course of 315 days, 0.93 million atoms of Carbon 14 decayed to 916341 radioactive atoms. Estimate the probability that on a given day, 36 radioactive atoms decayed.
4) Assume that 37 (sigma = 37) is the population's standard deviation. It is intended to calculate the population average (µ) for this population. An interval with a maximum size of 5 (half size 2.5) and a 92% confidence level will be produced for this purpose. What number of samples are…
0.389, 0.511 >
Describe the shape of the given histogram. Select the correct answer below
1. Per-item cost of overestimating demand: $3.28 2. Per-item cost of underestimating demand: $14.23 3. Probability that the value of demand is lower than the optimal order quantity: 0.50 4. Optimal order quantity for uniform distribution: 470 items 5. Optimal order quantity for normal…
A distribution of values is normal with a mean of 227.2 and a standard deviation of 50.6. Find P37, which is the score separating the bottom 37% from the top 63%. P37 =
David scored a 65% on the exam and claims that he did better than at least 50% of the other students. After all, 65% is higher than 50%. Do you agree with him?
Heights of females are known to follow a Normal distribution with mean 64.5 inches and standard deviation 2.8 inches. Based on this information, answer the following questions. Find the maximum height that would put a female in the bottom 5% of all female heights, first by hand and then using…
A recent survey shows that 60% of the households in a large area own 1 car. 9 households are selected at random. (a) What is the probability that at most 7 of these households having 1 car? (6 marks) (b) What are the expected number and variance of households owning 1 car? (4 marks) (c) Most…
.In a study to understand employee preferences for workplace training methods, a survey found that 83% of the employees from the tech department preferred online training modules, while 75% of the employees from the marketing department favored the same. Each department had 100 respondents in…
The correct hypothesis that Mifrin is testing is: Null Hypothesis (H0): The mean time to relief for headache sufferers is greater than or equal to 3.5 minutes. Alternative Hypothesis (H1): The mean time to relief for headache sufferers is less than 3.5 minutes.
A section of an exam contains four true/false questions. A completed exam is selected at random and the four answers are observed. Assume all outcomes are equally likely. (a) Describe the sample space. (b) Find the probability that all the answers are "true". (c) Find the probability that all…
Model 1: (B) Use a simple linear regression of y = Price against X = AGE. State your NULL & ALTERNATE hypotheses (for each of these models) and your conclusions. Use α = .05 throughout this project (for each of these models). Decide how much the value of a car is lowered if it is one year…
A random sample of 23 patients has an average time for a kidney stone to pass of 2.5 days and a standard deviation of 0.63. Find the 91% confidence interval for the population mean and round all values to two decimals.
1. A-F Suppose we observe x_(1),x_(2),dots,x_(n) as independent and identically distributed as an exponen- tial random variable with f_(lambda )(x_(1))=lambda e^(lambda _(1)), and recall that E(x_(1))=(1)/(lambda ) and Var(x_(1))=(1)/(lambda ^(2)). (a) Write down the conditional density…
For the given data set representing the runs scored by two players in the last 10 matches, conduct the following analysis: egin{center} egin{tabular}{|c|c|c|c|c|c|c|c|c|c|c|} hline Match & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 9 & 10 \ hline Player A & 8 & 42 & 56 & 68 & 91 & 123 & 12 & 46 & 57 &…
Problem 4 (15 points) A call center receives 25 callers per minute on average. On average, a caller spends 1 minute on hold and 3.5 minutes talking to a service representative. On average, how many callers are "in" the call center (meaning that they are either on hold or talking to a service…
A Box-and-Whisker plot for the sample data set "Some Data" is shown, along with some additional statistics about the data. ar{x} =48.9 s=16.5,n=83 a. Find each of the following: Range: Median: Interquartile Range: b. What proportion of the data falls between the values of 37 and 71 ? (Enter…
13.15 Suppose Y is a continuous random variable having the density function fY(y)=1/π(1+y^2) for y∈R. (a) Verify that this is a valid density function. (Hint: Use symmetry, and remember trig substitution!) (b) Try to find the expected value of Y and show that it does not exist. 13.17 Suppose…
Nico will perform a hypothesis test at the alpha =0.01 level. Evie will perform the same test at the alpha =0.10 level. Part: 0 / 2 Part 1 of 2 (a) If H_(0) is true, who has a greater probability of making a Type I error? If H_(0) is true, | has a greater probability of making a Type I…
From a binomial probability distribution with n=20, select any probability of success p between 0.1 and 0.4 from a cumulative binomial probability table and use the table to find: a) P(x>=5) and b) P(4<=x<=10) c) P(x=6) 3.From a binomial probability distribution with n-20, select any…
As N becomes smaller, x (standard error) becomes stays the same. larger. closer in value to the mean. smaller. As N becomes smaller,X (standard error) becomes stays the same. larger. closer in value to the mean. smaller.
5.5-2. Let x be N(75,64). Using the same set of axes, sketch the graphs of the probability density functions of (a) x. (b) ar{x} , the mean of a random sample of size 16 from this distribution. (c) ar{x} , the mean of a random sample of size 36 from this distribution. 5.5-2. Let X be N(75,…
(d) If an order comes in for 10,000 steel rods, how many rods should the plant manager expect to manufacture if the order states that all rods must be between 27.9 cm and 28.1 cm?
Predicting expected revenue in dollars from a new customer, predicting sales revenue for a store, and predicting the length of stay in a hospital based on patient characteristics and medical conditions are examples of: a. Applicability of statistical techniques b. Applicability of…
For a normal distribution, exactly 89.97% of the z-score values are less than z = +1.50. True False
For this exercise, you will be navigating problem 10.13, Sexually Transmitted Disease, and problem 10.15, Ophthalmology, both on page 439 of the textbook. Here are a few helpful hints: For problem 10.13, construct a 2x2 contingency table of the data presented in the problem, then use a…
Solution: There are 8 (= 2 * 2 * 2) different cases, and the results are provided in the following table: | n | σ | 1-α | 1-α/2 | Zα/2 | σZα/2 √n | Confidence Interval | |-----|-----|-----|-------|------|----------|----------------------| | 10 | 0.01| 95% | 97.5% | 1.96 | 0.062 | (Xn…
Question 3 10pts A test was given to a group of students. The grades and gender are summarized in the chart below. If one student is chosen at random from those who took the test, find the probability that the student was male given they got a B. Write your answer as a decimal rounded to 4…
A student-faculty government committee of 4 people is to be formed from 21 student volunteers and 4 faculty volunteers. Find the probability that the committee will consist of the following, assuming the selection is made at random: A student-faculty government committee of 4 people is to be…
accompanying table. able[[,y(x,y),0,5,10,15],[,0,0.01,0.06,0.02,0.10,,,,,],[x,5,0.04,0.13,0.20,0.10,,,,,],[,10,0.01,0.15,0.17,0.01,,,,,]] (a) Compute the covariance for x and Y. (Round your answer to two decimal places.) Cov(x,Y)= (b) Compute ho for x and Y. (Round your answer to two decimal…
Suppose the manufacturer can employ a temporary employee to work more hours in step S2 at $10 for each additional hour. Is it worthwhile? How much would the total cost change for one more hour? Group of answer choices It is worthwhile as it could save $7.5 for each additional hour. It is not…
1024,256,64,16,4, (Type an integer or a simplified fraction.) Determine the most probable next term in the list of numbers 1024,256,64,16,4 1024256,64,16,4, (Type an integer or a simplified fraction.)
Test whether β(0) is significant at α = 0.05. The p-value is 0.05. H0: We conclude that the intercept of the regression significantly contributes to the annual maintenance cost. What is the interpretation of β(0)? Is the interpretation of β(0) reasonable? The interpretation is because β(0) has…
Click the icon to view the table. Percentage of scores GREATER than 122 is ◻%. (Round to two decimal places as needed.) Data table Standard Scores and Percentiles for a Normal Distribution Full data set able[[Standard score,%, able[[Standard],[score]],% Use the normal distribution of IQ…
c. mu_{(D)}>0 d. mu_{(D)<0 b. What is the alternative hypothesis? a. mu_{(D)}=0 b. mu_{(D)} eq0 c. mu_{(D)}>0 d. mu_{(D)}<0 Data: able[[Action Movie,Horror Movie],[11,12],[15,18],[16,17],[20,21],[24,22],[25,23],[26,29],[13,16],[18,21],[19,23],[20,15],[24,25]] c. What are your degrees of…
Based on your sample, you will conduct a hypothesis test to test two of the claims of the above article. Using the same Excel sheet as last week, answer the following in the "Week 6" tab: Claim: The average age of online students is 32 years old. Can you prove it is not? - What is the null…
Hi, I will help you with this. Let's start by correcting the errors in the text provided: Original: (a) Show that P(x<=n)=1-(1-p)^(n),ninZ^(+). (b) Deduce an expression for p=0.2n The random variable X has a geometric distribution with parameter p (a) Show that P(Xn)=1-(l-p)", neZ*. [3 marks]…
A small-business owner collected data from a random sample on the amount spent on an order and the time it takes to fill the order. The resulting data were used to complete a linear regression analysis of time, in minutes, versus amount spent, in dollars, and to create the following residual…
Dr. Rankin was interested in examining the impact of mindfulness meditation on reducing stress levels. After collecting data, she calculated the mean and standard deviation for reduction in stress scores among participants. Further, she wanted to compare the collected data to a known population…
Can you also show a pixture od the answrs in excel Compute the correlation between motivation and GPA. Test for the significance of the correlation coefficient at the .05 level using a two-tailed test. True or false? The more highly you are motivated, the more you will study. Which did you…
Question 2: ***DISCRETE MATH*** You are celebrating the end of the semester at your local pub. Being an intrepid computer scientist, you run a randomized algorithm to determine what you order. Every time the wait staff asks you your order, you roll 2 fair, six-sided dice and add them together…
A machine that fills beverage cans is to be replaced by a new system if the variance in the new machine has a variance of less than 0.01 when filling up 12 ounces of beverage in a can. In order to test the variance in the new machine 10 cans are filled and the amount of liquid in each can is…
Calculate the minimum building airflow for the well-shielded, two-story residence. Building data: Living space area = 2400 ft^2 Number of occupants = 5 Number of stories above grade = 2 Location in Houston, TX egin{table}[h] centering egin{tabular}{|c|c|c|c|c|c|} hline Zone & Number of…
This data is from a sample. Calculate the mean, range, standard deviation, and coefficient of variation. Round answers to two decimal places. able[[x This data is from a sample. Calculate the mean, range, standard deviation, and coefficient of variation. Round answers to two decimal places. X…
Assume the random variable x is normally distributed with mean μ = 50 and standard deviation σ = 7. Compute the probability. Be sure to draw a normal curve with the area corresponding to the probability shaded. OA. P(34 < X < 57) probability shaded. 0 Which of the following normal curves…
(a) The transition probability of a Markov chain with threc states {0,1,2} is P=([(3)/(4),(1)/(4),0],[(1)/(4),(1)/(2),(1)/(4)],[0,(3)/(4),(1)/(4)]) and the initial state distribution of the chain is P{x_(0)=i}=(1)/(3),i=0,1,2. Find (i) P{x_(2)=2} and (ii)…
Show all steps with answers please. QUESTION 22 For the following data: Group 1 Group 2 8 3 5 4 6 9 4 8 2 7 Calculate the following: a.Ss 1 b.SS 2 c.SS D d. e.s
A journal on mental health presented the results of a survey of over 6,000 clients of a state's mental health department. One of the many variables measured for each mental health patient was frequency of social interaction (on a 5-point scale, where 1 = very infrequently, 3 = occasionally, and…
a. Using the results in column (1), what is the expected change in price of building a 500-square-foot addition to a house? Construct a 95% confidence interval for the percentage change in price. b. Comparing columns (2) and (3), is it better to use Size or ln(Size) to explain house prices?
Chapter 8: (8.2) Confidence Intervals (Mean) - sigma unknown - Homework Score: 13.8/16 9/10 answered Progress saved sqrt(0)◯,vdots (0)/(2) pts Details Assume that a sample is used to estimate a population mean mu . Find the 98% confidence interval for a sample of size 35 with a mean of 26.8…
You wish to test the following claim (HA) at a significance level of α = 0.10. Ho: P1 = P2 HA: P1 > P2 You obtain 506 successes in a sample of size n1 = 720 from the first population. You obtain 361 successes in a sample of size n2 = 543 from the second population. What is the test statistic…
Use the Standard Normal Table or technology to find the z-score that corresponds to the cumulative area 0.0071 . z= (Type an integer or decimal rounded to two decimal places as needed.) Use the Standard Normal Table or technology to find the z-score that corresponds to the cumulative area…
[20 points] Consider the problem of predicting how well a student does in her second year of college/university, given how well she did in her first year. Specifically, let x be equal to the number of "A" grades (including A-. A and A+ grades) that a student receives in their first year of…
It is known that the number of misprints on a textbook page follows a Poisson distribution with a mean of 0.004 . Which of the following is the standard deviation of the number of misprints on a textbook of 475 pages? A. 0.004 B. 0.063 C. 1.9 (e^(-lambda )*lambda ^(x))/(x!) D. 1.378 X! A.…
Which of the following come out to about the same number when they are calculated with the same set of data? a. mean absolute deviation and variance b. mean absolute deviation and standard deviation c. variance and standard deviation 10. Which of the following come out to about the same number…
Exercise 7. Let xi be a real random variable whose c.d.f is F_(xi ). Find the c.d.f's of the following random variables in terms of F_(xi ) : zeta =|xi |. eta =ln(xi ) (assuming xi >0 ). heta =exp(xi ). 1 Suppose that xi is normally distributed with mean mu and variance sigma ^(2), i.e its…
A psychologist wanted to know whether there were visual spatial differences in the development of children if they are raised within a hunter-gather culture. A sample of n=36 children between the ages of 6 - 8 years old were recruited and a cognitive test scores 145 . The general mu =120 with a…
Question 2 (2 points) You'd like to be able to compare scores on the SMART with another standard aptitude test -- the ACT. In order to do this, we need to transform the SMART distribution so that it has the same mean and standard deviation as the ACT. ACT has a mean of 20 and standard deviation…
You are playing a game with a friend. You roll a die. If you roll a 3 , you win $15. If you lose, she collects $3. Is this a fair game? Explain. 7. You are playing a game with a friend. You roll a die. If you roll a 3, you win $15. If you lose, she collects $3. Is this a fair game?Explain.
Calculate the Sensitivity of a screening test having these results: Of those with the disease: 1000 test positive, 100 test negative; of those without the disease 250 test positive and 2500 test negative. Answer as a percentage. Calculate the Sensitivity of a screening test having these…
An SRS of 50 students were selected from a population of 200 students. The population has a parameter equal to 0.85. What is the center and spread of the sampling distribution? In this example when calculating spread of sampling distribution do you divide by sample or by…
Pickmeup is a drive-through coffee house that has room for 3 cars in the driveway. The line cannot exceed 3 cars, and because they are exclusively drive-through, customers may be turned away. In the morning, the arrival rate is 40 cars per hour (Poisson distributed) and with the servers working…
Sample standard deviation S2 = 6. In performing the pooled-variance t-test, how many degrees of freedom are there? There are __ degrees of freedom. (Simplify your answer.) Assume there is a sample of n = 5, with the sample mean X1 = 35 and a sample standard deviation of S1 = 3, and there is an…
(1 point) Match the correlation coefficients with their scatterplots. Select the letter of the scatterplot below which corresponds to the correlation coefficient. (Click on image for a larger view.) r=0.22 r=0.89 r=-0.74 r=-0.49 4.1 Scatter Diagrams Linear Correlation: Problem 5 Previous…
According to a Pew Research Center project, the main sources of news information for Americans are television and:
Provided below are summary statistics for independent simple random samples from two populations. Use the pooled t-test and the pooled t-interval procedure to conduct the required hypothesis test and obtain the specified confidence interval. ar{x} _(1)=17,s_(1)=6,n_(1)=19,ar{x}…
Part 6 of 9 Points: 0 of 1 The regression model hat(Y)_(i) = -349.3 + 1.7513x_(1i) - 0.0912x_(2i) predicts standby hours based on total staff present, x_(1i), and remote hours, x_(2i), for week i. The data from which the model was constructed are provided. Use the regression model and the data…
What would be quantitative analysis be for the mathematical results of warning and actions Production 000000 Week1 Week2 Estimated Lost Production Week3 Week4 Mean Weeks Week5 UCL Week 6 LCL Week7 Week8 Control Chart LCL Mean UCL
Compute the probability of x successes, using the binomial formula. a. n=6,x=3,p=0.03 b. n=4,x=2,p=0.18 c. n=5,x=3,p=0.63 Compute the probability of X successes,using the binomial formula. a.n=6,X=3,p=0.03 b.n=4,X=2,p=0.18 c.n=5,X=3p=0.63
Age discrimination: The following table presents the numbers of employees, by age group, who were promoted, or not promoted, in a sample drawn from a certain industry during the past year. [ egin{array}{|c|c|c|c|c|} hline ext{Age} & ext{} & ext{Under 30} & ext{30-39} & ext{40-49} &…
In a board game, a player rolls a 6 sided dice with the numbers 1 to 6 on it, and spins a spinner with the letters A to D on it. a. Create an outcome table to list all the possible outcomes. (1 mark) able[[,1,2,3,4,5,6],[A,,,,,,],[B,,,,,,],[C,,,,,,],[D,,,,,,]] N a.s with the letters A to D on…
2. You believe that the mean number of cats present in Guney Campus Manzara at any given moment is 4. Please answer each part below separately. a. You have counted the number of cats 9 times, and ar{x} turned out to be 3.3. If a frequent visitor of the place told you that sigma^2 for the…
You are a financial analyst who is evaluating the quarterly sales of two different companies. Based on the following data (numbers are in millions of dollars), construct an ANOVA table, state your null and alternate hypotheses, and indicate your conclusion based on a 0.05 level of…
Which of the following statements best describes current conservative attitudes about immigration? a) There is no division of opinion among conservatives because almost all conservatives strongly oppose immigration. b) There is no division of opinion among conservatives because almost all…
The data below are the temperatures on randomly chosen days during a summer class and the number of absences on those days. Calculate the correlation coefficient, r. able[[Temperatur, x,75,88,94,93,91,101,78,103,83],[Number of absences, y,5,9,12,12,10,17,6,17,7]] A. 0.819 B. 0.890 C. 0.980 D.…
Data is collected to determine if the proportion of occupants in cars who wear seat belts (p_(1)) is less than the proportion of occupants in pickup trucks who wear seat belts (p_(2)). Which hypotheses are appropriate to test…
b. What is the probability that weekly demand will be greater than the number of available radios? The probability is 0.10 c. What is the expected weekly profit from the sale of the alarm clock radio? (Remember: There are only five clock radios available) The expected weekly profit is $◻.…
Calculate A_(x) in terms of u,y, and z. Problem 4 Given: (10 points) (i) Ax:n= u, (iii) Ax+n = z. Calculate Ar in terms of u, u, and z
For arbitrary random variables x and Y, prove that Cov(x,E[Y|x])=Cov(x,Y). 4.For arbitrary random variables X and Y,prove that Cov(X,E[Y|X])=Cov(X,Y)
QUESTION #9 (15 marks) How can engaged couples improve their chances of a happy marriage? A study selected 19 couples and tested the quality of their communication skills 3 months before marriage. Leading up to the marriage, the couples were then asked to participate in a communication skills…
*DISCRETE MATH* Question 4: We roll a fair red die and a fair blue die. We define the following random variables. R={(1 if the red die is even ):} 0 otherwise B={(1 if the blue die is 3 or less ):} 0 otherwise S={(1 if the sum of both dice is even ):} 0 otherwise It should be clear that R…
A two-echelon distribution network [one central distribution center (x) and two stores A and B)] has 120 units of a product at the central distribution center, and no additional receipts are expected at the central distribution center for at least a week. The on-hand inventory and…
10. Events A and B are such that P(A) = , P(B) = and P(A|B) = 11. (a) Find P(A ∩ B). (b) Find P(A ∪ B). (c) State with a reason whether or not events A and B are independent.
Exercise 10-11 CAS Exam 3 Spring 2005 #22 Drivers are classified as "good" or "bad." Results of the classification are assumed to be binomially distributed, with probability of being "good" equal to p. A sample consists of 100 drivers. Determine the critical value for testing the hypothesis…
Question 3 [1.0 pt: a (0.4 pts), b (0.6 pts)] Imagine we want to estimate the population mean of some variable Y (e.g. earnings). To keep things simple, think of taking a random sample of size two from this population. Write Y_(1) and Y_(2) for the two values you get from the sample. Here are…
Example 10-6 A quality inspector counted the number of defective parts a prototype machine made in samples of 100 taken every hour during a 20-hour period. a) Using the following data, construct a Three Sigma control chart for the sample proportion of defectives. b) Is the machine producing a…
Determine the area under the standard normal curve that lies between (a) Z=-1.35 and Z=1.35, (b) Z=-0.77 and Z=0, and (c) Z=-1.65 and Z=1.18. (a) The area that lies between Z=-1.35 and Z=1.35 is (Round to four decimal places as needed.) Round to four decimal places as needed.
A researcher conducted a left-tailed hypothesis test on the amount of caffeinated coffee drunk by coffee drinkers, with HA: μ < 23. They found the p-value to be 0.0799, using a significance level of α = 0.10. Select the 2 best answers: Select all that apply: - Fail to reject the Null…
Data: 198,188,208,205,210,198,189,183,190,195,176, 189,174,155,180 c. What are your degrees of freedom? d. What is the critical t value ( t table)? e. What is the obtained t value? f. What can you conclude? a. Reject the null hypothesis. American men at a southwestern college are significantly…
A.) Show that k=2/9 B.) Find the mean and standard deviation of x
Bill keeps the phone numbers for his seven closest friends(three men and four women) in his digital phone memory. In how many ways can he list them if (a) men are listed before women? (b) men are all listed together? (c) no two men are listed next to each other? 7. Bill keeps the phone numbers…
(5 points) The five-number summary for a set of data is as follows: Minimum: 25, Q1: 32, Median: 40, Q3: 47, Maximum: 60. If there are 250 values in the data set, about how many values are between 32 and 47? (Hint: Think about percentiles.) 5.5 pointsThe five-number summary for a set of data…
Fifty people participated in a medical study. Each participant contributed two eye measurements (labeled y ). Thus the study had a total of 100 measurements on y from the 50 participants. It is suspected that the two eye measurements from the same person are correlated. Without sufficient…
Confidence Intervals for One Proportion Worksheet Math 181 Statistics Helgeson Due at the start of class on Monday, April 1, 2024. On the Sections 2.1, 2.2 & 2.3 (Theory-Based Hypothesis Testing for a Single Variable) lecture notes, we used simulation and also a theory-based hypothesis test…
6 and 7) Question 2 of 15 This question: 1 point(s) possible Survey results indicate that σ = 15.8 books. A 99% confidence level requires subjects. (Round up to the nearest whole number as needed.) 6 and 7) Question 2 of 15 This question: 1 point(s) possible People were polled on how many books…
In a class with 40 students, 18 are boys, and 22 are girls. In the last exam, 2 boys and 2 girls achieved grade A. Calculate the probability of, in a random selection of two people, a boy or a student with grade A being selected. (3)/(4) (1)/(2) (4)/(40) (18)/(40) In a class with 40 students,…
Coefficient to assess the normality of the data. Click here to view the data table. Click here to view the table of critical values for normal probability plots. Data Table (b) The correlation is (Round to three decimal places as needed.) able[[Index,…
What is the probability of randomly selecting a respondent who is more conservative than "moderate"? What is the probability of randomly selecting a respondent who is either "slightly liberal" or "slightly conservative"? What is the probability of randomly selecting two respondents, the first…
Question 8 (1 point) Assume we want to run a hypothesis test for the population mean. If we were to find a p-value of 0.0431 , what decision should we make at the 0.05 level of significance? Don't reject the null. There is NOT evidence of the alternate. Accept the null. There is evidence of the…
Two dice are rolled, find the probability of on which the sum of the dice is less than 18. Two dice are rolled,find the probability of on which the sum of the dice is less than 18
To understand the numbers in this article, you need to understand "median wage". To figure out the median wage, you would list the wages in order from smallest to largest and then find the wage where half of the people make less and half of the people make more (50th percentile). This is the…
A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 237.1-cm and a standard deviation of 2.2-cm. Find the probability that the length of a randomly selected steel rod is greater than 240-cm. P(x>240-cm)= Enter your answer as a number accurate to…
Use the following data set (data file provided) to determine if a significant statistical difference between 3 groups in the following experiment. Please perform a One-Way ANOVA analysis (ANOVA Single Factor), and then interpret the data. In this experiment, subjects perform a 1-hour…
. From the statistical and stochastic test scores of 100 students, 40 samples were randomly selected, namely: able[[,53*,57*,62*,65,71*,74,83],[42*,54*,58,62,65,72*,75,86],[45*,55,58,64,68*,73,76,86],[49*,56*,58,64,70*,73,78,86],[50*,56*,60,64*,71*,74,82,87]] a. What is the lowest value that…
Bag A contains 2 red and 3 green balls. a) Two balls are chosen at random from the bag without replacement. Find the probability that 2 red balls are chosen. Bag B contains 4 red and n green balls b) Two balls are chosen without replacement from this bag. If the probability that two red balls…
Hi need Help ASAP, thank you! Anil is a candidate in an election. He received 40% of the votes. A random sample of 120 voters is chosen. Use an approximation to find the probability that, of the 120 voters, between 36 and 54 inclusive voted for Anil. Anil is a candidate in an election. He…
For the following information, determine whether a normal sampling distribution can be used, where p is the population proportion, alpha is the level of significance, hat(p) is the sample proportion, and n is the sample size. If it can be used, test the claim. Claim: p>=0.34;alpha =0.04.…
A politician claims that the mean salary for managers in his state is more than the national mean, $82,000. Assume the population is normally distributed and the population standard deviation is $7600. The salaries (in dollars) for a random sample of 30 managers in the state are listed. At α =…
A population of values has a normal distribution with mu = 240 and sigma = 71.8. You intend to draw a random sample of size n = 42. Find the probability that a single randomly selected value is less than 217.8. P(X < 217.8) = Find the probability that a sample of size n = 42 is randomly…
QUESTION 5 What happens to the length of the confidence interval as the standard error increases? The length of the confidence interval increases. The length of the confidence interval decreases. There is no change to the length of the confidence interval as long as alpha doesn't…
Given two independent samples, conduct a hypothesis test for the difference between the means. Consider a significance level of 0.05 . Use a confidence interval to decide whether to reject or not H_(0). H_(0):mu _(1)-mu _(2)=0 H_(1):mu _(1)-mu _(2)!=0 able[[Sample (1),Sample…
Question 4 Use the data in the following table, which lists survey results from undergraduate school students. Assume that subjects are randomly selected from those included in the table. If one of the students is randomly selected, find the probability that the selected student took online…
Ignore the questions that talk about RStudio (the questions with yellow highlighter) In the class of 2021, where scores for all students who took the SAT during high school. For Problem Set 6, you have 524 students. 1. What is the probability that a student has a SAT Math score lower than…
Fill out the combined proxy column 7.7 + (3.7 x Production Cost) Fill out the combined proxy column 7.7 + (3.7 x Production Cost) + (7.6 x Promotion Costs) + (0.8 x Book Sales) Box Office Receipts Production Costs Promotion Costs Book Sales Combined Proxy 85.1 8.5 5.1 4.7 81.6 106.3 12.9 5.8…
A survey collected data about the number of TVs that a family of four has in their household. The number of TVs is distributed as shown. Find the mean, variance, and standard deviation for the distribution. (Write answers rounded to one decimal place. Use the letter A or B to answer the…
Find the EZ and Var(Z) for the Z∼(mu ,sigma ). Find the EZ and Var(Z) for the Z~(u,o)
The hypothesis is that there is no difference in the proportion of all men and women in the city who intended to vote for Trump. The null hypothesis is that the proportion of men intending to vote for Trump is equal to the proportion of women intending to vote for Trump. The alternative…
computers as factory C.The probability that a computer produced by factory A is defective is 0.04, the probability that a computer produced by factory B is defective is 0.033, and the probability that a computer produced by factory C is defective is 0.046. What is the probability it came from…
What is the correct sequence of events (oldest to youngest) for this diagram? Enter the letters with no spaces or commas. For example: ABCD What is the correct sequence of events (oldest to youngest) for this diagram? Enter the letters with no spaces or commas. For example: ABCD B c D A I
Problem 1 If x_(1),...,x_(n) are iid from Unif(- heta , heta ) RVs. Find a consistent estimator of heta . Problem 1 If X.....Xn are iid from Unif(-0,9) RVs. Find a consistent estimator of 0.
The analysis of results from a leaf transmutation experiment (turning a leaf into a petal) is summarized in the table by type of transformation completed. Let A denote the event that a leaf completes the color transformation, and Let B denote the event that a leaf doesn't complete the textural…
Model 1 (A) Use a simple linear regression of y= Price against x= Odometer. State your NULL & ALTERNATE hypotheses (for each of these models) and your conclusions. Use α = 0.05 throughout this project (for each of these models). Plot the graph of price against odometer and add it to your…
Consider a game where 3 contestants each roll an n-sided die. On the first play of the game, each will roll a 3-sided die. On the next play of the game, each contestant will roll a 4-sided die. On each subsequent play, the number of sides on the die will increase by 1. a) Suppose on any play…
It's New Year's Day, 2023, and thanks to a long night out with friends, you are off to a slow start. Little Cottonwood Canyon is closed due to extreme avalanche conditions. You decide to check out the weather at Snowbasin. Snowbasin's Boardwalk observing site provides snow interval observations…
ch babies first crawl. [ egin{array}{|c|c|c|c|c|c|c|} hline 49 & 31 & 46 & 34 & 39 & 27 \ hline 49 & 39 & 54 & 26 & 37 & 28 \ hline end{array} ] Chi-Square Distribution Critical Values Table Click the icon to view the table of critical values of the chi-square distribution Select the correct…
SUMMARY OF FORMULAS FORMULA 6.1 FORMULA 6.2 Confidence interval for a sample mean (large sample, population standard deviation known): Confidence interval for a sample mean (large sample, population standard deviation unknown): Confidence interval for a sample proportion (large…
For a standard normal distribution, find: P(0.39 < z < 2.18)
The size of fish is important to commercial fishing. A study conducted in 2102 found that the length of Atlantic cod caught in nets in Karlskrona, Sweden is normally distributed with mean 49cm and standard deviation 4cm. Source: Condition indices of Atlantic cod (Gadus morhua) biased by…
5.2 Some Probability Rules Compound Events: Problem 16 Results for this submission At least one of the answers above is NOT correct. ( 1 point) In a certain community, 35% of the families own a dog, and 20% of the families that own a dog also own a cat. It is also known that 29% of all the…
Let {Y_(t)} be a stationary process with model equation Y_(t)=24+(2)/(3)Y_(t-1)+epsi lon_(t)+(1)/(2)epsi lon_(t-1),tinZ, where {epsi lon_(t)} is a white noise process such that the epsi lon_(t) are NID(0,sigma ^(2)). Let y_(1),y_(2),dots,y_(T) denote an observed realization of the process up to…
Suppose we estimate the regression y = β0 + β1x1 + β2x2 + u, using 282 observations. We then conduct several hypothesis tests with a significance level α = 0.05: F-test for goodness of fit, two Breusch-Pagan tests, and two White tests (using the predicted dependent variable in a cubic…
PRACTICE ANOTH The method of tree-ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution. able[[1194,1292,1292,1250,1268,1316,1275,1317,1275]] (a) Use a calculator with mean and standard…
Suppose X1, . . . , Xn are iid RVs with mean lambda and variance lambda . Find a large sample 95% confidence interval for √lambda .Problem 2 If x_(1),dots,x_(n) are iid Unif(0, heta ). Find a large sample 95% confi- dence interval for heta . Problem 2 If X1,..:,Xn are iid Unif(0,0). Find a…
The Normal Approximation to the Binomial Distribution Consider the probability that less than 14 out of 157 flights will be delayed. Assume the probability that a given flight will be delayed is 12%. Approximate the probability using the normal distribution. Round your answer to three decimal…
2 Question 2 able[[Age,17,21,64,3 r.,86,41,26,58,46],[Income,10,500,55,367,45,678,37,985,65,786,126,121,97,360,67,587,278,941]] Sean is a 27 year old making an income of $85,500. Is he making above the predicted value of 2 Question 2 Age 17 21 64 86 41 26 58 46 Income 10,500 55,367 45,678…
The sample mean and standard deviation from a random sample of 29 observations from a normal population were computed as ar{x} = 28 and s = 15. Calculate the t statistic of the test required to determine whether there is enough evidence to infer at the 4% significance level that the population…
Interviewed in a random order. (a) What is the probability that x of the top five candidates are interviewed on the first day? h(x; 7, 5, 11) h(N; 4, 5, 11) h(N; 7, 11, 5) h(x; 4, 11, 5) (b) How many of the top five candidates can be expected to be interviewed on the first day? (Round your…
Assume that a person spins the pointer and is awarded the amount indicated by the pointer. Determine the person's expectation. A. $4.50 B. $3.25 C. $6.50 D. $4.25 To determine the person's expectation awarded the amount indicated by the pointer, assume that a person spins the pointer and is…
Question 6 7 pts According to admissions data from the Massachusetts Institute of Technology (MIT), 4.1% of freshman applicants are admitted to the college. If 33,240 individuals applied to MIT last year, find the probability that at least 1,400 of them were admitted. Round your answer to 4…
We have this population: {2,3,4,5,6}. The mean is 4 , but it is ignored by the investigator. He still insists to estimate the mu from samples. We have three possible estimators which are: Estimator 1: sample mean Estimator 2: sample median Estimator 3: average of the largest and the smallest…
Problem 3 Suppose x_(1),dots,x_(n) are iid RVs with mean lambda and variance lambda . Find a large sample 95% confidence interval for sqrt(lambda ). Problem 3 Suppose X..... Xn are iid RVs with mean X and variance A. Find a large sample 95% confidence interval for VX
Two men A and B toss a coin in turn until a head occurs. The man who tosses a head wins. a) Find the probability that A wins. b) If the coin is biased and head occurs with probability p, what is the probability that B wins? c) Let X be the number of trials required to produce the 1st success,…
In Box I are bulbs that are guaranteed to last one year, selling for R3 each. In Box II are bulbs with no guarantee selling for R1 each. The store owner explains that Box II came to him unlabeled. He knows that all of the bulbs in Box II are of one type and declares that it is equally likely…
A single sample of 9 data points is taken. If you perform a right-tailed one sample hypothesis test for the means with an alpha value of 0.05, what is your degree of freedom value and what is the t-critical value for this test.
The correct answer is: c. X2 < 29.1413
When it comes to people snoring while sleeping, many will use an anti-snoring device. About 70% of users lose 10 pounds in two weeks. Create a hypothesis test to evaluate this claim. Represent it as a null hypothesis and an alternative hypothesis.
Let G be a set containing 5 elements a, b, c, d, and e. Under a binary operation called the multiplication, we know that ab = c and cd = dc = e. Show that G is not a group under this operation
Find the critical values chi Subscript Upper R Superscript 2 and chi Subscript Upper L Superscript 2 for the given confidence level c and sample size n. c equals 0.80, n equals 16 Question content area bottom Part 1 chi Subscript Upper R Superscript 2 equals    enter your response here (Round…
9. (Normally distributed quality characteristic) Suppose a quality characteristic of a product is normally distributed with mean 18 and standard deviation 1.5 . The specification limits furnished by the customer are (15,21). Determine what percentage of the product meets the specifications…
Find the regression line for data points X | Y 85 |80 90|40 100| 30 102| 26 105| 20
Consider a continuous random variable X with probability density function f(x) = c · x(1 − x)^5, 0 ≤ x ≤ 1,
A random sample of 12 cats was collected to study the lifespan with the data listed below. Find the 87% confidence interval for the population mean. Round all values to two decimals. 13, 15, 18, 16, 8, 12, 13, 15, 8, 13, 12, 11. Provide your answer below.
a) A simple random sample of size 20 has mean x = 3.19. The population standard deviation is a = 1.51. The population is approximately normally distributed. Construct a 90% confidence interval for the population mean.
What is the empirical probablity of obtaining 9 heads in 10 tosses
Project for MA235 STEP 1: You are doing research on a topic of interest. Decide what you want to research. To complete this research you will need to determine a question that has a numerical response. Identify a random sample of n = 15 men and n = 15 women. NOTE: Though it is suggested, you do…
I need some help interpreting some data. I have some of the answers but I am having a hard time finishing this. 1. Describe the method used and why you used the method. We will apply this by sending out emails of our survey to 300 random students at BYU-Idaho. We chose this method because we…
Ciorstaidh's investment has a 20% chance of producing a 14% return, a 50% chance of producing a 9% return, and a 30% chance of producing a -2% return. What is its standard deviation? a. 4.46% b. 5.81% c. 6.00% d. 0.36%
Find the regresion line: X: 40, 60, 80, 100, 120, 140, 160. Y:0, 0.2, 0.4, 0.6, 0.8, 1, 1.2
Let be a random variable with a standard normal distribution. Find the indicated probabilities. a) P (z<-1.32)= ____________ c) To the left of z=72, _____________________ b) P(z<=-2.15)=__________________ d) P(z>=2.17) =_____________ e) P(z>=-1.50)= ___________
Describe a forensically relevant scenario that might involve a binomial random variable and thus require the use of the Binomial distribution
The probability distribution that can be described by just one parameter is the _____ distribution. a. uniform b. continuous c. normal d. exponential
If lambda is set at 34, what is the probability of 29 or more but no more than 46?
a) Suppose there is population with 4 customers. The interested categorical variable is the proportion of customers that prefer your brand. In this population, one of them prefers your brand, and the others do not. (i) Find the population proportion (pi ) if customers that prefer your brand. …
I need the R codes and the answers to the question. Using R Studio, load wage1 data from the wooldridge package. Estimate the model log(wage) = B0 + B1educ + B2exper + B3tenure + u. Interpret the coefficients and the model overall. Calculate the VIF values. Do you think the multicollinearity is…
As a part of the study of the development of thymus glands, researchers weighed the glands of chick embryos after 14 days of incubation. The thymus weights (mg) were as follows: 29.6, 30.1, 24.7, 21.5, 33.4, 29.8, 41.5, 32.9, 33.7, 28.0, 34.6, 44.9, 37.8, 23.4, 27.1, 41.5, 43.2 a) Calculate…
According to a survey, the mean student loan at graduation is $50,000. Suppose that student loans are normally distributed with a standard deviation of $5,000. A graduate with a student loan is selected at random. Find the following probabilities. (Round your answers to four decimal…
Here's a list of units of toys sold by T&K over a 12-month period during 2015. Were the sales of 31,456 for one month in 2016 significantly different from monthly sales in 2015? Units Sold in 2015 January 34,518 February 29,540 March 34,889 April 26,764 May 31,429 June 29,962 July …
A bag contains 3 boxes of sour candy and 5 boxes of chocolates. (a) If you draw one box at random, what is the probability that it is a box of sour candy? (b) A friend asks for a box of chocolates and removes it from the bag. What is the probability now of drawing a box of sour candy at random?
Using the interquartile range which of the following are outliers select all correct answers 43 4448 6470
A random sample of 5 patients has an average time for a kidney stone pass of 10.2 days and a sample standard deviation of 0.39. Find the 89% confidence interval for the population mean. Round all values to two decimals. Provide your answer below.
Why is an understanding of different scales of measurement(nominal, ordinal, interval and ration) important for the purpose of questionnaire design?
A random sample of six nursing students has an average amount of sleep in 24 hours of 7.6 hours and a sample standard deviation of 3.78. Find the 79% confidence interval for the population mean. Round all values to two decimals. Provide your answers below.
7 Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 5.8-in and a standard deviation of 0.9-in. In what range would you…
Part 3: Hypothesis Test: (2 points) I had recently conducted an experiment: the moment the fuel light in my car goes on, I will find a closest gas station to fill up the tank. But the number of gallons needed to fill up are all over the place. This makes me wonder how we can verify it is…
A firm specializing in agricultural products wants to conduct a test-market offering for one of its new products. A random sample of 800 potential customers is chosen to receive advertising material describing the new product, and it was determined that 600 out of the 800 people made a…
If, in a sample of n = 30 selected from a normal population, X̄ = 52 and S = 12, what are the critical values of t if the level of significance, α = 0.05, the null hypothesis, H0, is μ = 50, and the alternative hypothesis, H1, is μ ≠50?
A distribution of values is normal with a mean of 124.5 and a standard deviation of 4.7. Find the probability that a randomly selected value is greater than 138.1. P(X > 138.1) =
What is the probability that a randomly selected baby born this year weighed 3000 to 3999 grams or was postterm
For a standard normal distribution, find: P(z > -2.72) Express the probability as a decimal rounded to 4 decimal places.
In a bag there are 10 balls. There are 4 red balls and the remaining balls are green balls removed at random and the color noted, the ball is replaced. Then a second ball is removed at random and the color is noted. P(1 red ball and 1 green ball picked)= 4/10*6/10=6/25 P(2 red balls picked…
5.40. A multiple-choice test consits of eight questions and three answers to each question (of which only one is correct). If a student answers each question by rolling a balanced die and checking the first answer if he gets a 1 or 2, the second answer if he gets a 3 or 4, and the third answer…
If two events A and B are such that P(A) = 2/3, P(B|A) = 1/2 and P(A') = . Use this information to answer questions 5-11. 5. Find P(A|B'). 6. Calculate P(A|B). 7. Calculate P(A'|B). 8. Calculate P(A'|B'). 9. Find P(A ∩ B'). 10. Find P(A' ∩ B). 11. Find P(A' ∩ B'). 12. Suppose that A and B are…
An agent sells life insurance policies to 6 equally aged, healthy people. According to recent data, the probability of a person living in these conditions for 25 years or more is 2/5. Calculate the probability that after 25 years: A.) Find the probability that all 6 people are still living. (10…
Tree heights: Cherry trees in a certain orchard have heights that are normally distributed with mean mu =117 inches and standard deviation sigma =1 4 What is the probability that a randomly chosen tree is between 90 and 102 inches tall?
Read the following question about survey data collection and provide a brief response Bruce, a biology professor, is interested in opinions about the summer research program at his college and decides to conduct a survey of students who've participated. He makes the survey available online and…
A researcher conducted a study with 12 patients and found the sample mean of the kidney stone passing to be 6.4 days with a sample standard deviation of 3.4. They want to test if this is greater than the regional average of 9 days at a significance level of a=0.25. What is the p-value for this…
Select the 99% CI of mean differences using the following: df = 9 SE = 4.2 Mean difference = 5.5 a. (-10.16, 15.16) b. (2.21, 6.31) c. (-8.15, 19.15) d. (-7.22, 18.14)
Find the area under the standard normal curve that lies between z = -1.28 and z = -0.29.
suppose over the course of one hour, a player can expect to be dealt about 40 hands. how much should a player expect to win or lose over the course of three hours
For the following questions, do not compute numerically. (a) Compute the number arrangements (i.e. permutations) of 52 cards in a standard deck, in which red cards (hearts and diamonds, not necessarily in that order) appear before black cards (spades and clubs, not necessarily in that…
A random sample of 420 observations produced a sample proportion equal to 0.32. Find the critical and observed values of z for the following test of hypotheses using α = 0.1. Ho: p = 0.30 versus H1: p ≠0.30. Round your answers to two decimal places. Zcritical left = Zcritical right = Zobserved
A) For a binomial distribution with a sample size equal to 4 and a probability of a success equal to 0.18, what is the probability that the sample will contain exactly two successes? Use the binomial formula to determine the probability. Question content area bottom Part 1 The probability that…
d¯ = -1.000 sd = 4.899
A driver's reaction time to visual stimulus, R, is normally distributed with a mean of 0.4 seconds and a standard deviation of 0.05 seconds.a. Create the probability density function of R.b. Determine the mean, variance, and standard deviation of R.c. Plot the probability density function of…
A university has admitted a group of n students with varying skill levels. To better accommodate the students, the university has decided to create classes tailored to the skill levels. A placement examination will return a skill level that will be used to group the students, where levels[i]…
The following products are included in the current State Farm Health product portfolio EXCEPT: -Short Term Disability Insurance - Long Term Disability Insurance -Long Term Care Insurance -Medicare Supplement Insurance The monthly benefit range for the Short Term Disability Insurance policy…
Students are given a multiple choice exam with 4 questions. The random variable X = the number answers correct. Possible values = {0, 1, 2, 3, 4}. From past data, 10% of students get zero correct answers, 10% get exactly one correct answer, 20% get two correct, and 40% get three correct. It can…
Problem 4. (15 points) Let X denote the lifetime of a radio, in years, manufactured by a certain company. Assume that the lifetime is exponentially distributed, with a mean of 5 years. (a) (5 points) What is the probability p that a radio lasts more than 15 years? (b) (5 points) What is the…
Customers arrive at a pharmacy to be served according to a Poisson process with rate lambda . Knowing that q <= n, m <= n, and q + m <= n, and that 0 < t1 < t2 < t: (a) (3 points) If n customers arrive in the interval [0, t], what is the probability that exactly q customers have arrived in the…
In a laboratory where the temperature is being kept at the range of 20 degrees and 25 degrees it is found that over the course of 315 days, 0.93 million atoms of Carbon 14 decayed to 916341 radioactive atoms. Estimate the probability that on a given day, 36 radioactive atoms decayed.
4) Assume that 37 (sigma = 37) is the population's standard deviation. It is intended to calculate the population average (µ) for this population. An interval with a maximum size of 5 (half size 2.5) and a 92% confidence level will be produced for this purpose. What number of samples are…
0.389, 0.511 >
Describe the shape of the given histogram. Select the correct answer below
1. Per-item cost of overestimating demand: $3.28 2. Per-item cost of underestimating demand: $14.23 3. Probability that the value of demand is lower than the optimal order quantity: 0.50 4. Optimal order quantity for uniform distribution: 470 items 5. Optimal order quantity for normal…
A distribution of values is normal with a mean of 227.2 and a standard deviation of 50.6. Find P37, which is the score separating the bottom 37% from the top 63%. P37 =
David scored a 65% on the exam and claims that he did better than at least 50% of the other students. After all, 65% is higher than 50%. Do you agree with him?
Heights of females are known to follow a Normal distribution with mean 64.5 inches and standard deviation 2.8 inches. Based on this information, answer the following questions. Find the maximum height that would put a female in the bottom 5% of all female heights, first by hand and then using…
A recent survey shows that 60% of the households in a large area own 1 car. 9 households are selected at random. (a) What is the probability that at most 7 of these households having 1 car? (6 marks) (b) What are the expected number and variance of households owning 1 car? (4 marks) (c) Most…
.In a study to understand employee preferences for workplace training methods, a survey found that 83% of the employees from the tech department preferred online training modules, while 75% of the employees from the marketing department favored the same. Each department had 100 respondents in…
The correct hypothesis that Mifrin is testing is: Null Hypothesis (H0): The mean time to relief for headache sufferers is greater than or equal to 3.5 minutes. Alternative Hypothesis (H1): The mean time to relief for headache sufferers is less than 3.5 minutes.
A section of an exam contains four true/false questions. A completed exam is selected at random and the four answers are observed. Assume all outcomes are equally likely. (a) Describe the sample space. (b) Find the probability that all the answers are "true". (c) Find the probability that all…
Model 1: (B) Use a simple linear regression of y = Price against X = AGE. State your NULL & ALTERNATE hypotheses (for each of these models) and your conclusions. Use α = .05 throughout this project (for each of these models). Decide how much the value of a car is lowered if it is one year…
A random sample of 23 patients has an average time for a kidney stone to pass of 2.5 days and a standard deviation of 0.63. Find the 91% confidence interval for the population mean and round all values to two decimals.
1
...
16
17
18
19
20
...
22